OpenStudy (annie96):

Question 5.5. Solve, using the substitution method. y = 3x + 5 4x – y = 5 (Points : 1) (10, 35) (15, 10) There are an infinite number of solutions. There is no solution.

jigglypuff314 (jigglypuff314):

y = 3x + 5 <-- plug this one into the other one 4x – y = 5 -> 4x - (3x + 5) = 5 distribute the negative 4x - 3x - 5 = 5 add like terms add 5 to both sides x = (something) <-- (x, y) ^ which option has what you got for x there

OpenStudy (annie96):

(10,35)

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (annie96):

Thank you.

jigglypuff314 (jigglypuff314):

you're welcome :)